OpenStudy (anonymous):

GIVING MEDAL AND BE A FAN!!!! if f(x)=2x, g(x)=2x-1, and h(x)=radical of x, find (f o g o h) (9)

OpenStudy (anonymous):

oops i meant what is \(h(9)=\sqrt{9}\) ?

OpenStudy (anonymous):

(f o g o h) = f(g(h(x)))=f(g(sqrt(x)))=f(2*sqrt(x)-1)=2*(2*sqrt(x)-1) when x is 9, 2*(2*sqrt(9)-1)=10.

OpenStudy (anonymous):

woooooaaaahhh

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

\[h(9)=\sqrt{9}=3\] \[g(3)=2\times 3-1=5\] \[f(5)=2\times 5=10\]